Estimated demand for cars is normally distributed with a standard deviation of 20 and a mean of 120. What is the probability that:


(i)More than 160 cars will be needed. (4 marks)

(ii)Less than 90 cars will be needed. (4 marks)

(iii)More than 100 but less than 130 cars will be needed.(4 marks)


1
Expert's answer
2021-02-01T12:06:40-0500

Mean μ = 120

SD σ = 20

(i)

"P(X>160) = 1 -P(X<130) \\

= 1 -P(\frac{X-μ}{σ} < \frac{160-120}{20} ) \\

= 1 -P(Z<2) \\

= 1 -0.9772 \\

= 0.0228"

(ii)

"P(X<90) = P(\frac{X-μ}{σ} < \frac{90-120}{20}) \\

= P(Z<-1.5) \\

= 0.0668"

(iii)

"P(100<X<130) = P(\frac{100-120}{20}<\frac{X-μ}{σ}< \frac{130-120}{20}) \\

= P(-1<Z<0.5) \\

= P(Z<0.5) -P(Z<-1) \\

= 0.6914 -0.1586 \\

= 0.5328"